Evaluate: 32 + (8 - 2) · 4 - 6/3
Answer by Student A(4)

You start by solving what is in the parentheses : 8-2 = 6 Then you do multiplication or division depending on which comes first as you move from left to right. Here multiplication is first so : 6x4= 24. Then the question becomes 32 + 24 - 6/3 . 6/3 = 2 => 32+24-2=54. So your answer is 54.